The volume of a cuboid is 3000cm³, surface area 1300cm² and diagonal 5√29cm. What is the ratio of its length, breadth and height ?

Grade:10

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

H = volume/ surface area = 30/13 cm

L^2 + b^2 = ( 25*29) - (900/169)

L*b = 1300

From here we can calculate required

Regards